A gratting of width 2inches is ruled with 15000 lines per inch. Find the smallest wavelength separation that can be resolved in second order at a mean wavelength of 500nm
1
Expert's answer
2018-09-17T15:31:08-0400
R=λ/∆λ=mN.

m=2,N=15000∙2=30000
So,

R=mN=2∙30000=60000.

We have

λ/∆λ=mN → ∆λ=λ/mN=(500∙〖10〗^(-9))/60000=0.0083∙〖10〗^(-9) m=8.3 pm.